OpenStudy (anonymous):

^3√32 simplify into radical form

OpenStudy (anonymous):

\[\sqrt[3]{32}\] Can you make prime factors for 32??

OpenStudy (anonymous):

8 and 4

OpenStudy (anonymous):

Yes: Do you know what is cube root of 8??

OpenStudy (anonymous):

See, do you know even how to take cube root??

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

2??

OpenStudy (anonymous):

the cube root of 8 is 2 @waterineyes right?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

\[\sqrt[3]{32} = \sqrt[3]{8 \times 4} \implies \color{green}{2 \sqrt[3]{4}}\]